numerator of fraction increased by 12% & denominator decreased by 2% the value of fraction becomes 6/7. thus the original fraction is
Answer by fcabanski(1391)

If x is increased by 12% it is multiplied by 1.12:x + .12x = 1.12x


If y is decreased by 2%, it is multiplied by .98: y -.02y = .98y


1.12x/.98y = 6/7


You can multiply either side by 1 without changing its value. Multiply the left side by 1 in the form of 100/100 to remove the decimals.


112x/98y = 6/7


Multiply both sides by 98/112.


x/y = 6/7 * 98/112 = 6/7 * 7/8 = 6/8 = 3/4.